bambi235504:

Which of the following evaluates to 9? A.10c1 B.9c8 C.12c10 D.9c9

TheSmartOne:

Hello @bambi235504 ! Welcome to QuestionCove!

TheSmartOne:

Do you know the difference between permutations and combinations and the formulas for them?

TheSmartOne:

Well, since you're gone now. I'll just drop the formulas: \(\Large nPr = \frac{n!}{(n-r)!}\) \(\Large nCr = \frac{n!}{r!(n-r)!}\)